The performance curve of a pump is represented by the following function:

h=266.7-.6667q2

where h is head in feet, and q is discharge in cfs.

If three of these pumps were used in series, what would be the head at a discharge of 12.4 cfs?
Answer #1

Head and discharge relations

h= (266.7-0.6667*q^2) ( ie head will decrease as discharge increases)

when the pumps are in series,the discharge will remain constant and head will be additive of the 3 pumps.

for discharge, q = 12.4 cfs,

head,h = 266.7-0.6667*(12.4^2) = 164.19 feet

so total head at discharge 12.4 cfs when 3 pumps are in sereis = 3*164.19 = 492.56 feet (ans)